OpenStudy (jackfrostt3482):

Please help asap I will give medal and fan. How do you multiply a fraction 1/3 for example with pi 3.14??

OpenStudy (jackfrostt3482):

@zepdrix @.Sam. @jabez177 @AloneS @Awolflover1 @imqwerty @King.Void. @Ashleyisakitty @ShadowLegendX @inkyvoyd @Elsa213 @Jamierox4ev3r @563blackghost @Hayhayz @Jaynator495 @iwanttogotostanford @tshimp0629 @GIitter @desmarie @ethane @HanaMalik @Unavailabilityy @bonnieisflash1.0 @psirockin2 @ryandean123 @goodfaze123 @babybritt125 @Christopher1448 @HipPopGirl

jabez177 (jabez177):

\[3\frac{ 14 }{ 100 }~or~\frac{ 314 }{ 100 }~=~\frac{ 22 }{ 7 }\]

OpenStudy (jackfrostt3482):

wha..?

jabez177 (jabez177):

Then you'll do \[\frac{ 1 }{ 3 }\times \frac{ 22 }{ 7 }\]because 22/7 is about equal to pi.

OpenStudy (jackfrostt3482):

Ok :) thank you!